HUBBLE SPACE TELESCOPE DAILY REPORT #2122 PERIOD COVERED: 0000Z (UTC) 05/12/98 - 0000Z (UTC) 05/13/98 Daily Status Report as of 133/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.4 Completed Six Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.5 Completed Five Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed NIC/2 7862 (A Search for Near Infrared H-2 Emission in Active Elliptical Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of active elliptical radio galaxy 3C278 to search for infrared H-2 region emissions. The observations were completed as scheduled, and no anomalies were reported. 1.7 Completed NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.8 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed STIS/CCD/WF/PC-2/NIC/2 7388 (Black Holes and the Centers of Galaxies) The Space Telescope Imaging Spectrograph (CCD), the WF/PC-2, and the Near Infrared Camera (Camera Two) were used to continue our survey of the center of nearby galaxy NGC-3608. The major goal is to survey an unbiased set of galaxies with a potentially wide range of black hole masses. The sample has been carefully chosen to optimize leverage on the character of a proposed correlation of black hole mass with galaxy mass. The observations were completed as scheduled with no anomalies. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Operations Notes: The STIS MCE reset at 132/155851Z while in an SAA contour and while the high voltage was off. HSTAR 6472 was written. The STIS error counter was reset at 132/1912Z using ROP NS-12. Then, before the MCE-1 (MAMA1) was recovered, another MCE reset occurred. This was at 133/060118Z, and during another SAA event. HSTAR 6474 was written. The MAMA1 will be recovered via the SMS at 133/1818Z. Science operations were not affected by either of these events. The SSR correctable EDAC counts were cleared once per ROP SR-1. It was reported today by HSTAR 6473 that at 122/200700Z the section 5 current dropped from approximately 11 amps to 7.6 amps. This value continued to decay until it reached 4 amps over the next six minutes, with two transient increases to approximately 10 amps during the decay. At 122/201340Z, the section 5 current returned to approximately 11 amps (the normal value, with K51 open, since the occurrence of the section 4 to section 5 short on day 064). These described changes were also reflected in the battery charge current. There is no immediate health and safety issue with the vehicle. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH